what is the equation for __ is what percent ___? for example: 7 is what percent of 28?
Answer by CharStar(110)

You can use the technique of solving proportions when working on percent problems. Use is/of = rate/100. % goes over 100.
7/28=x/100 = Cross Multiply
7 times 100 = 700
28 times X = 28x
x = 25
To check you anser 28 X 25 = 700 and 7 X 100 = 700
Enjoy your day